OpenStudy (anonymous):

Which ordered pair (p, r) is a solution to the given system of linear equations?

OpenStudy (anonymous):

5p-3r=1 8p+6r=4 answer choices (-1,-2) (-2,-1) (2/9,1/3) (1/2,2/9) neep help asap

OpenStudy (anonymous):

do you know how to apple elimination or substitution method?

OpenStudy (anonymous):

*apply

OpenStudy (anonymous):

make them equal to zero?

OpenStudy (anonymous):

using the elimination method means you have to eliminate one variable in order to solve for the value of the other variable. Substitution method, you have to rearrange one of the equations in terms of a variable and substitute it to the next equation in order to solve for another variable. Now which method do you want to use?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

i think the answer is A though?

OpenStudy (anonymous):

have you tried solving it? did you get A?

OpenStudy (anonymous):

yes, i got A

OpenStudy (anonymous):

i don't think it's right because i got a different one. can you show me your solution?

OpenStudy (anonymous):

what about this one? Veda solves the following system of linear equations by elimination. What is the value of x of the solution? 6+4x-2y=0 -3-7y=10x -20 -1 1 20

OpenStudy (anonymous):

how do i take a picture? i thought it was either D or A

OpenStudy (anonymous):

Its D! I plugged the answers in

OpenStudy (anonymous):

first rearrange the equations: 4x-2y=-6 10x+7y=-3 let's eliminate y by multiplying eqn 1 by 7 and eqn to by 2 7(4x-2y=-6) 28x-14y=-42 2(10x+7y=-3) 20x+14y=-6 now, add the two equations you'll get: 48x=-48, so x=-1

OpenStudy (anonymous):

wow, awesome okay! i need help with 2 more

OpenStudy (anonymous):

this time you have to do it:) i'll check how you do it

OpenStudy (anonymous):

André solves the following system of equations by elimination. 3a+6b=12 -3a+6b=-12 what is the resulting equation? 12b=0 6a=0 -6a=0 -12b=-24

OpenStudy (anonymous):

okay i can try, but im on a timed test with 15 minutes remaining and one question left

OpenStudy (anonymous):

oh wait, i can do that one

OpenStudy (anonymous):

sorry but we are just here to guide you for you to get the answer and not to provide answer because that's considered cheating

OpenStudy (anonymous):

Which statement correctly explains how Andre should find a solution to the following system of linear equations using elimination? 8s-9t=15 7s-2t=12 Multiply the first equation by –7 and the second equation by 8, and then subtract. Multiply the first equation by 7 and the second equation by 8, and then subtract. Multiply the first equation by 2 and the second equation by 9, and then add. Multiply the first equation by –2 and the second equation by 9, and then subtract. stuck!

OpenStudy (anonymous):

okay well ive been plugging in for each problem and im not getting an answer for this one

OpenStudy (anonymous):

what's the purpose of elimination? as i've mentioned before: using the elimination method means you have to eliminate one variable in order to solve for the value of the other variable. so what should you do to eliminate one variable in the equations?

OpenStudy (anonymous):

eliminate the s or t?

OpenStudy (anonymous):

right, so what should you do to the equations to eliminate one of them?

OpenStudy (anonymous):

add in 0 for one of them?

OpenStudy (anonymous):

nope. try applying the choices on the equations. If a choice eliminates one of the variables then that would be the right one, if not, try another one.

OpenStudy (anonymous):

ive plugged them all in and i keep getting 2 for most of the answers.

OpenStudy (anonymous):

ill keep trying

OpenStudy (anonymous):

don't PLUG the values, APPLY them. say for example, the last choice: Multiply the first equation by –2 and the second equation by 9, and then subtract. so first equation multiply by -2 it will be -2(8s-9t=15) = ? then 2nd equation multiply by 9 it will be 9(7s-2t=12) =? then subtract the two equations and see if one variable is eliminated or not.

OpenStudy (anonymous):

thank you! i got an 80 on the test:)

OpenStudy (anonymous):

can i still use yall for help?

OpenStudy (anonymous):

if we will help you, please try to help yourself also because we are not here to provide you with answers. it's against our Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i understand, i didnt expect answeres. thank you

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

:) welcome to openstudy

OpenStudy (anonymous):

At a farmer’s market, a basket of strawberries costs $20 and a basket of peaches costs $18. On Saturday, $430 worth of strawberries and peaches were sold, and a total of 23 baskets were sold. If x represents baskets of strawberries and y represents baskets of peaches, how many baskets of peaches were sold on Saturday? so so far i have written down the different variables